A hydrometer whose mass is 25g is made up of a bulb of diameter 2cm and height 5cm and a stem of diameter 1cm and height 10cm.
Determine the range of specific gravities that the hydrometer can be employed to measure?
1
Expert's answer
2021-08-30T01:55:03-0400

"Specific \space gravities = \frac{Density \space of \space object }{Density \space of \space water}\\
Specific \space gravities = \frac{\frac{Mass}{Volume} }{Density \space of \space water}\\
Specific \space gravities = \frac{\frac{0.025}{ \pi *0.02^2*0.05}+\frac{0.025}{ \pi *0.01^2*0.1} }{1000}\\
\frac{\frac{0.025}{\pi 0.02^2\cdot \:0.05}+\frac{0.025}{\pi 0.01^2\cdot \:0.1}}{1000}\\
=\frac{\frac{0.025}{0.00002\pi }+\frac{0.025}{0.00001\pi }}{1000}\\
=1.19366"